OpenStudy (anonymous):

f(x)=x^2-7x+9 Find the x and y intercept

OpenStudy (anonymous):

idk if this will help but f(x + h) − f(x) h For f(x) = 2x + 3 we have: f(x + h) − f(x) h = [2(x + h) + 3] − [2x + 3] h = 2x + 2h + 3 − 2x − 3 h = ✚2✚x + 2h + 3✁ − ✚2✚x − 3✁ h = 2h h = 2

OpenStudy (alekos):

The y intercept is when x=0 So just find f(0)

OpenStudy (alekos):

The x intercept requires that we try and factorise the quadratic

OpenStudy (campbell_st):

remember f(x) and y are interchangeable. for the y intercept let x = 0 so you have \[y = 0^2 -7 \times 0 + 9\] for the x-intecepts you have to set y = 0 and solve for x. \[0 = x^2 - 7x + 9\] and looking at this equation, you'll need the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you have a = 1, b = -7 and c = 9 just substitute and calculate the values. hope it helps

OpenStudy (alekos):

x^2 -7x +9 = (x+a)^2+b

OpenStudy (alekos):

did you work out the intercepts?